Problem B
Show that 3 · 4n + 51 is divisible by 3 and 9 for all positive integers n.
1
Expert's answer
2020-07-29T12:50:17-0400

For any integer n ≥ 1, let "P_n" be the statement that "3\cdot4^n+51" is divisible by 9.


Base case. The statement "P_1" says that "3\cdot4^1+51=63" is divisible by 9, which is true.

Inductive Step. Fix k ≥ 1, and suppose that "P_k" holds, that is, "3\cdot4^k+51" is divisible by 9,


"3\cdot4^k+51=9m, m\in \Z, m\geq7"

It remains to show that "P_{k+1}" holds, that is, "3\cdot4^{k+1}+51" is divisible by 9,


"3\cdot4^{k+1}+51=4\cdot(3\cdot4^k)+51=""=4\cdot(3\cdot4^k+51)-4\cdot51+51="

"=4\cdot9m-3\cdot51=9(4m-17),m\in \Z, m\geq7"

Therefore "P_{k+1}" holds.

Thus, by the principle of mathematical induction, for all n ≥ 1, "P_n" holds.

Therefore "3\cdot4^n+51" is divisible by 9 for all positive integers n.

Therefore "3\cdot4^n+51" is divisible by 3 for all positive integers n.